# appReferences in the Appendix

### v2

@@ 2.2.2 Subgradient-based Algorithm
Thus, this process can be interpreted as a projected stochastic subgradient descent method for solving the optimization problem (2). The method reduces computational costs by requiring only a single pass through the data and eliminates the need for solving LPs.
@@ 2.3 Performance Metrics
We evaluate the algorithms using a bi-objective performance measure that involves the optimality gap and constraint violation. The optimality gap measures the difference between the objective value of the algorithm’s output and that of the true optimal solution, while the constraint violation measures the degree to which the algorithm’s output fails to meet the given constraints.
We denote the order-request matrix by $\bm{A}=[\bm{a}_{1},...,\bm{a}_{T}]^{\top}$, the offline optimal solution to problem (PLP) by $\bm{x}^{*}=(x_{1}^{*},...,x_{T}^{*})^{\top}$, and the online optimal solution by $\bm{x}=(x_{1},...,x_{T})^{\top}$. Then, we define the regret $r(\bm{x})$ and resource violation $v(\bm{x})$ as:
$r(\bm{x})$ $\coloneqq\max_{\mathbf{A}\mathbf{x}\leq\bm{b},\mathbf{0}\leq\mathbf{x}\leq\mathbf{1}}\leavevmode\nobreak\ \leavevmode\nobreak\ \sum_{t=1}^{T}r_{t}x_{t}^{*}-r_{t}x_{t},$ (8)
@@
Therefore, we define the following bi-objectives for evaluating the total regret and worst-case total regret of the algorithms we propose:
@@
where $\mathcal{P}$ is the unknown distribution of $\{(r_{t},\bm{a}_{t})\}_{t=1}^{T}$ and $\Xi$ denotes the family of distribution satisfying some regularity assumptions specified later.
This metric is commonly used for first-order OLP algorithms in [7, 15] and is also aligned with the literature on the online convex optimization with constraints in [18, 24]. For LP-based methods, [16] use stopping time to bound the regret. Equation (10) integrates the analysis of subgradient methods and LP-based methods into a single performance metric, enabling more consistent analysis across online learning scenarios. This integrated method promotes more balanced resource consumption over the decision horizon. This flexibility makes our framework well-suited for scenarios without prior knowledge of the constraint support, thereby expanding its applicability to a broad range of online learning and resource allocation problems.
@@ 2.4 Assumptions and Auxiliary Results
We propose the following assumptions regarding the stochastic inputs used throughout the paper. These minimal assumptions are standard in the online learning literature [16, 13, 23]. In particular, we require the input data to be bounded, follow a linear growth, and be non-degenerate.
Assumption 1 (Boundedness).
@@
    There exist constants $\bar{r},\bar{a}>0$ such that $|r_{t}|\leq\bar{r}$ and $\lVert\bm{a}_{t}\rVert_{\infty}\leq\bar{a}$ almost surely for $t=1,...,T$.
@@
    The average resource capacity $\bm{d}=\bm{b}/T$ satisfies $d_{i}\in[\underline{d},\bar{d}]$ for some $\bar{d}>\underline{d}>0$ for any $i=1,...,m$.
In this assumption, (a) states that $\{(r_{t},\bm{a}_{t})\}_{t=1}^{T}$ are independent of each other, but we allow dependencies between their components. Part (b) introduces the bounds $\bar{r},\bar{a}$ solely for analytical purposes; they are not used within the algorithm itself. This is a minimal requirement on $(r_{t},\bm{a}_{t})$ compared to previous work [23]. Part (c) requires the average resource capacity to grow linearly with $T$, ensuring that a constant fraction of the $x_{t}$ values can be set to 1 in the optimal solutions. Consequently, the number of fulfillable orders is proportional to $T$, facilitating a stable service level over time.
Assumption 2 (Uniform Non-degeneracy).
@@
    The second moment matrix $\mathbb{E}[\bm{a}\bm{a}^{\top}]$ is positive definite with minimum eigenvalue $\lambda$.
@@
    There exists constants $\mu,\nu$ such that for any $(r,\bm{a})\sim\mathcal{P}$,
    $\nu|\bm{a}^{\top}(\bm{p}-\bm{p}^{*})|\leq|\mathbb{P}(r\geq\bm{a}^{\top}\bm{p}\mid\bm{a})-\mathbb{P}(r\geq\bm{a}^{\top}\bm{p}^{*}\mid\bm{a})|\leq\mu|\bm{a}^{\top}(\bm{p}-\bm{p}^{*})|$
    holds for all $\bm{p}\in\mathcal{V}_{\bm{p}}\coloneqq\{\bm{p}\in\mathbb{R}^{m}:\bm{p}\geq\bm{0},\lVert\bm{p}\rVert\leq\frac{\bar{r}}{\underline{d}}\}$ and $\bm{d}\in\mathcal{V}_{\bm{d}}\coloneqq[\underline{d},\bar{d}]^{m}$.
@@
    The optimal solution $\bm{p}^{*}$ satisfies $p_{i}^{*}=0$ if and only if $d_{i}-\mathbb{E}[a_{i}\mathbb{I}(r>\bm{a}^{\top}\bm{p}^{*})]>0$ for all $\bm{d}\in\mathcal{V}_{\bm{d}}$ and $i=1,...,m$.
In this assumption, part (b) ensures that the cumulative distribution of the reward given the resource consumption request $r|\bm{a}$ is continuous and exhibits a stable growth rate. Part (c) requires strict complementarity for the optimal solutions of the stochastic program in (3), which is a non-degeneracy condition for both the primal and dual LP.
@@
$I_{B}$ =$=\{i:d_{i}-\mathbb{E}[a_{i}\mathbb{I}(r>\bm{a}^{\top}\bm{p}^{*})=0]\},$
$I_{N}$ =$=\{i:d_{i}-\mathbb{E}[a_{i}\mathbb{I}(r>\bm{a}^{\top}\bm{p}^{*})>0]\}.$ (11)
Here, $I_{B}$ denotes the set of binding resources—those that are limited and require careful allocation—while $I_{N}$ denotes the set of non-binding resources, which are unlimited and can fulfill any order’s request. By Assumption 2(c), these sets are complements, as $I_{B}\cap I_{N}=\emptyset$ and $I_{B}\cup I_{N}=\{1,...,m\}$.
Let $\Xi$ denote the family of distributions satisfying Assumptions 1 and 2. We propose the following lemma.
Lemma 2.1 (Dimension Stability).
Under Assumptions 1 and 2, there exists a constant $\delta>0$ such that
[…]
This lemma is a consequence of Lemmas 12 and 13 in [16]. The existence of $\delta$ comes from the continuity of $f_{\bm{d}_{t}}(\bm{p})$. Thus, $\delta$ is only associated with stochastic program (7), and it is independent of $T$. Note that $\mathcal{D}\subset\mathcal{V}_{\bm{d}}$. We will analyze our feasible dual solutions derived from the online algorithm with $\bm{d}_{t}\in\mathcal{D}$.
Lemma 2.2 (Quadratic Regularity, Proposition 2 in [16]).
Under Assumptions 1 and 2, for any $\bm{p}\in\mathcal{V}_{\bm{p}}$,
Moreover, $\bm{p}^{*}$ is the unique optimal solution to (3).
3 Parallel Multi-Phase Online Learning and Decision-Making
In this section, we present our novel framework for parallel multi-phase online learning and decision-making. Our approach is motivated by the difficulties of existing online resource allocation methods, which involve either high regret or heavy computational costs. Specifically, the LP-based AHDL Algorithm 1 has a strong worst-case regret bound of $\mathcal{O}(\log T)$, but its high computational costs delay decision-making. Meanwhile, the first-order online Algorithm 2 updates decisions efficiently but suffers a high regret of $\mathcal{O}(\sqrt{T})$. By combining the strengths of these two methods, our new framework balances decision-making quality and computational efficiency, thereby overcoming the limitations of existing approaches.
@@ 3.1 Algorithm Design
    We maintain two parallel processes—one dedicated to online learning and the other to online decision-making. The online learning results are periodically sent to the decision-making path to guide subsequent updates.
@@
    In the online learning path, we employ an enhanced LP-based AHDL algorithm, which only re-solves the updated OLP problem according to a predefined frequency, significantly reducing the computational overhead.
@@
    In the online decision-making path, we apply the first-order method exclusively during the initial and final re-solving batches, where the learning rate is optimally tuned for these two intervals.
Figure 1: Parallel Multi-Phase Online Learning and Decision-Making Algorithm
Figure 1 illustrates our framework, which generates two parallel sequences of dual solutions: $\{\bm{p}_{t}^{D}\}_{t=1}^{T}$ from a first-order method and $\{\bm{p}_{t}^{L}\}_{t=1}^{T}$ from an LP-based approach. The algorithm proceeds in batches of length $f$. In the first batch, it uses a first-order method to iteratively update dual prices, thereby guiding the decision-making process. At the end of this batch, it applies the LP-based AHDL algorithm to obtain a refined set of dual prices, passing these updated values to the decision-making path. This LP-based re-solving occurs only once per batch, making the approach computationally efficient. Formally, we re-solve the OLP at every time $t$ satisfying:
$t\leq kf\quad\text{and}\quad t\bmod f=0,$ (12)
where $k=\left\lfloor\frac{T}{f}\right\rfloor$ represents the number of batches.
After the first batch, the algorithm does not employ the first-order method until the final batch. Instead, during intermediate batches, the decision-making path uses the most recent dual prices computed from the LP at the end of the previous batch. Only in the final batch does the algorithm restart the first-order updates to guide the final set of decisions. Algorithm 3 summarizes this approach.
Algorithm 3 Parallel Multi-Phase Online Learning and Decision-Making
Input: total resource $\bm{b}$, time horizon $T$, average resource $\bm{d}=\bm{b}/T$, initial dual price $\bm{p}_{1}=\bm{0}$, re-solving frequency $f$, and number of re-solving batches $k=\lfloor T/f\rfloor$.
for $t=$ $1$ to $T$  do
       Observe $(r_{t},\bm{a}_{t})$ and make decision $x_{t}$ based on rule (6). Update constraint for $i=1,...,m$:       remaining resource constraint $b_{it}=b_{i,t-1}-a_{it}x_{t}$       average resource capacity $d_{it}=\frac{b_{it}}{T-t}$ Update learning rate $\alpha_{t}=\mathcal{O}(1/t)$. if $t\leq f$ or $t\geq kf$ then
             Compute the subgradient and update the dual price $\bm{p}_{t+1}$:
$\bm{p}_{t+1}$ $=(\bm{p}_{t}-\alpha_{t}(\bm{d}-\bm{a}_{t}x_{t}))\vee\bm{0}$ (13)
@@
             Solve the online dual problem and update dual price $\bm{p}_{t+1}$ as
$\bm{p}_{t+1}=\operatornamewithlimits{arg\,min}_{\bm{p}\geq\bm{0}}\ \bm{d}_{t}^{\top}\bm{p}+\frac{1}{t}\sum_{j=1}^{t}(r_{j}-\bm{a}_{j}^{\top}\bm{p})^{+}$ (14)
@@
We achieve a strong synergy between an effective online learning method and an efficient decision-making procedure. This structure enables the algorithm to remain responsive to dynamic environments and adapt to the latest information. Our strategy of periodic re-solving greatly reduces the computational costs. Therefore, we establish a “wait-less” online decision-making framework in which each customer’s order is processed immediately without waiting for the completion of earlier requests.
@@ 3.2 Algorithm Analysis
We next present several lemmas that characterize the convergence of the online dual solutions $\{\bm{p}_{t}\}_{t=1}^{T}$ generated by Algorithm 3 toward the optimal solution $\bm{p}^{*}$ of the stochastic program (3). In particular, let $\bm{p}_{t}^{D}$ be the dual prices updated via the first-order rule (13), and let $\bm{p}_{t}^{L}$ be those obtained from the LP-based method (14) at time $t$. These lemmas will then be used to establish upper bounds on the worst-case regret of our algorithm.
Lemma 3.1 (Dual Price Boundedness).
Under Assumption 1, the online and offline optimal dual prices are bounded respectively by:
$\lVert\bm{p}^{*}\rVert\leq\frac{\bar{r}}{\underline{d}},\ \ \ \lVert\bm{p}_{t}\rVert\leq\bar{p}\coloneqq\max\left(\frac{\bar{r}}{\underline{d}-\delta},\ \frac{2\bar{r}+m(\bar{a}+\bar{d})^{2}}{\underline{d}}+m(\bar{a}+\bar{d})\right).$ (15)
This lemma establishes that the optimal dual prices remain bounded. Our algorithm maintains these bounds because if $\bm{p}_{t}$ grows large, the algorithm responds by accepting more orders, which in turn reduces $\bm{p}_{t+1}$. This self-correcting mechanism keeps the dual prices within these limits. Moreover, drawing on results of [15], we have that
$\lVert\bm{p}_{t}^{D}\rVert\leq\frac{2\bar{r}+m(\bar{a}+\bar{d})^{2}}{\underline{d}}+m(\bar{a}+\bar{d})$
and from [16], we have that
$\lVert\bm{p}_{t}^{L}\rVert\leq\frac{\bar{r}}{\underline{d}-\delta}.$
Consequently, we can prove the following convergence lemmas.
Lemma 3.2 (Dual Convergence of LP-based algorithm).
Under Assumptions 1 and 2, there exists a constant $C_{lp}>0$ depending on $\bar{r},\bar{a},\underline{d},m,\nu$, and $\lambda$ such that
$\mathbb{E}\left[\lVert\bm{p}_{t}^{L}-\bm{p}_{t}^{*}\rVert^{2}\right]\leq\frac{C_{lp}}{t}.$ (16)
In addition, the difference between $\bm{p}_{t}^{*}$ and $\bm{p}^{*}$ satisfies
This lemma establishes that the LP-based online AHDL algorithm [13] produces dual solutions $\bm{p}_{t}$ that converge to $\bm{p}_{t}^{*}$. This convergence highlights the stability of the online dual variables during the intermediate stages of decision-making and ensures a warm start for the first-order method in the last re-solving batch. Moreover, we bound the distance between $\bm{p}_{t}^{*}$ and $\bm{p}^{*}$ by relating their respective average resource capacities, $\bm{d}_{t}$ and $\bm{d}$, in the associated stochastic programs [16]. These results provide the foundation for analyzing $\lVert\bm{p}_{t}-\bm{p}^{*}\rVert$ in our final result.
Lemma 3.3 (Dual convergence of First-order algorithm).
Under Assumptions 1 and 2, if $\alpha_{t}<\nu\lambda$, the subgradient updates satisfy the following recursion rule:
$\mathbb{E}\left[\lVert\bm{p}_{t+1}^{D}-\bm{p}^{*}\rVert^{2}\right]\leq(1-\alpha_{t}\nu\lambda)\lVert\bm{p}_{t}^{D}-\bm{p}^{*}\rVert^{2}+\alpha_{t}^{2}m(\bar{a}+\bar{d})^{2}.$ (18)
$\mathbb{E}\left[\lVert\bm{p}_{t}^{D}-\bm{p}^{*}\rVert^{2}\right]\leq C_{fo}\left(\frac{1}{\alpha t}+\alpha\right).$ (19)
$\mathbb{E}\left[\lVert\bm{p}_{t}^{D}-\bm{p}^{*}\rVert^{2}\right]\leq\frac{C_{fo}}{t}.$ (20)
This lemma shows that the first-order method [8] guarantees the convergence of $\bm{p}_{t}$ to $\bm{p}^{*}$. It also highlights a key trade-off in choosing the learning rate $\alpha_{t}$, an aspect that will be central to our later optimality analysis. With this groundwork in place, we now move on to bounding the total regret of our algorithm.
Theorem 3.1 (Total Regret).
Under Assumptions 1 and 2, the total regret $\Delta_{T}$ of Algorithm 3 satisfies:
$\Delta_{T}\ \leq\ \mu\bar{a}^{2}\sum_{t=1}^{T}\mathbb{E}\left[\lVert\bm{p}_{t}-\bm{p}^{*}\rVert^{2}\right]+\lVert\bm{p}^{*}\rVert\cdot\mathbb{E}\left[\lVert(\bm{b}-\bm{Ax})^{B+}\rVert\right]+\mathbb{E}\left[\lVert(\bm{Ax}-\bm{b})^{+}\rVert\right].$ (21)
where $(\cdot)^{B+}$ indicates that we only consider the binding terms and project them onto the positive orthant.
We express the total regret as the sum of two key components: dual convergence and resource consumption. This bound demonstrates that the algorithm encourages (1) smooth and balanced resource utilization, especially for the binding terms, and (2) full resource consumption by the end of the time horizon. Building on Theorem 3.1, we next introduce a “spectrum theorem” for the algorithm’s total regret.
Theorem 3.2 (Spectrum Theorem).
Under Assumptions 1 and 2, the worst-case regret $\Delta_{T}$ of Algorithm 3 is bounded by
$\Delta_{T}\in\mathcal{O}\left(\sqrt{f}+\log\left(\frac{T}{f}\right)+f^{1/3}\right)=\mathcal{O}\left(\log\left(\frac{T}{f}\right)+\sqrt{f}\right),$ (22)
@@
Remark 3.1.
The Spectrum Theorem elucidates the trade-offs in total regret induced by our use of LP-based and first-order methods. As stated in Section 2, utilizing pure LP-based method at every step yields a regret of $\mathcal{O}(\log T)$, while applying pure first-order method continuously attains a regret of $\mathcal{O}(\sqrt{T})$. Our algorithm sets a re-solving frequency $f\in[1,T]$ and achieves the regret of $\mathcal{O}(\log(T/f)+\sqrt{f})$, which aligns with previous results with the extreme cases of $f=1$ and $f=T$. By selecting an appropriate re-solving frequency $f$, our algorithm leverages the strengths of both approaches. Specifically, first-order method gives us the regret of $\sqrt{f}$ for the first batch and $f^{1/3}$ for the last batch, and LP-based method produces the regret of $\log(T/f)$.
Remark 3.2.
We can obtain a tighter bound of
$\Delta_{T}\in\mathcal{O}(\log(T)+f^{1/3})$ (23)
if we have a warm start of the initial dual price for the first batch satisfying $\lVert\bm{p}_{0}-\bm{p}^{*}\rVert\leq f^{-1/3}$ or we use the LP-based method to learn at each time. Both these approaches can give us the improvement from $\sqrt{f}$ to $f^{1/3}$ for the first-batch regret.
@@
We prove this theorem using several novel techniques. As stated in Theorem 3.1, we develop a comprehensive framework that accommodates both the first-order and LP-based methods, allowing us to decompose the total regret into distinct components associated with each approach. Specifically:
  • For the first-order method, we relate the distance $\bm{Ax}-\bm{b}$ to the distance between dual prices.
  • For the LP-based method, as the algorithm progresses toward the end of the time horizon, if the current average resource consumption rate $\bm{d}_{t}$ exceeds the allowed limit, we manually set the dual prices to zero to accept all future orders. This approach enables us to adapt the concept of a “stopping time” within our new framework and derive a rigorous bound of regret.
Through this new framework, we integrate the LP-based and subgradient-based methods to achieve “wait-less” decision-making, where each customer’s order is processed immediately without delays, while also attaining the best-known logarithmic regret bound.

### v3

@@ 2.2 Algorithms for OLP
This process can be interpreted as a projected stochastic subgradient method for solving (1). It reduces computational cost by requiring only a single pass through the data and eliminates solving LPs explicitly. A restart strategy improves the first-order method to have $\mathcal{O}(T^{1/3})$ regret (Gao et al., 2024). Algorithm˜4 and Algorithm˜5 in Section˜A.3 summarizes this method.
@@ 2.3 Performance Metrics
We aim to design algorithms that optimize a bi-objective performance measure involving regret and constraint violation. The regret measures the difference between the objective value of the algorithm’s output and that of the true optimal solution, while the constraint violation measures the degree to which the algorithm’s output fails to meet the given constraints. We denote the offline optimal solution to (PLP) by $\bm{x}^{*}=(x_{1}^{*},\ldots,x_{T}^{*})$, and the online algorithm output by $\bm{x}=(x_{1},\ldots,x_{T})$. Then, we define the regret $r(\bm{x})$ and resource violation $v(\bm{x})$ as
$r(\bm{x})$ $\coloneqq\langle\bm{r},\bm{x}^{*}\rangle-\langle\bm{r},\bm{x}\rangle,$ (8)
@@
Therefore, we define the following bi-objectives for evaluating an algorithm’s worst-case performance:
@@
where $\Xi$ denotes a family of distributions satisfying regularity assumptions specified later.
This metric is commonly used for first-order OLP algorithms (Gao et al., 2023; Li et al., 2020) and is also aligned with the literature on online convex optimization with constraints (Mahdavi et al., 2012; Yu et al., 2017). Integrating $r(\bm{x})$ and $v(\bm{x})$ into a single performance measure promotes balanced resource consumption over the decision horizon. In Section˜3, we derive a decomposition of (10) that unifies the regret analysis for all OLP methods.
@@ 2.4 Assumptions and Auxiliary Results
We adopt the following assumptions regarding the stochastic inputs. These assumptions are standard in the online learning literature (Li & Ye, 2022; Jiang et al., 2022; Xu et al., 2024). In particular, we require the input data to be bounded, follow a linear growth, and be non-degenerate.
Assumption 2.1 (Boundedness).
@@
    There exist constants $\bar{r},\bar{a}>0$ such that $|r_{t}|\leq\bar{r}$ and $\lVert\bm{a}_{t}\rVert_{\infty}\leq\bar{a}$ almost surely for t=1,,Tt=1,...,T.
@@
    The average resource capacity $\bm{d}=\bm{b}/T$ satisfies $d_{i}\in[\underline{d},\bar{d}]$ for some $\bar{d}>\underline{d}>0$ for any i=1,,mi=1,...,m.
In this assumption, (a) states that $\{(r_{t},\bm{a}_{t})\}_{t=1}^{T}$ are independent of each other, but we allow dependencies between their components. Part (b) introduces the bounds $\bar{r},\bar{a}$ solely for analytical purposes. This is a minimal requirement on $(r_{t},\bm{a}_{t})$ compared to previous work (Agrawal et al., 2014; Li & Ye, 2022; Xu et al., 2024). Part (c) requires the average resource to grow linearly with $T$, ensuring that a constant fraction of $x_{t}$ values can be set to 1. Consequently, the number of fulfillable orders is proportional to $T$, facilitating a stable service level over time.
Assumption 2.2 (Uniform Non-degeneracy).
@@
    The second-moment matrix $\mathbb{E}[\bm{a}\bm{a}^{\top}]$ is positive definite with minimum eigenvalue $\lambda$.
@@
    There are constants $\mu,\nu$ such that for any $(r,\bm{a})\sim\mathcal{P}$,
    $\nu|\bm{a}^{\top}(\bm{p}-\bm{p}^{*})|$
    $\leq\$$\left|\mathbb{P}(r\geq\bm{a}^{\top}\bm{p}\mid\bm{a})-\mathbb{P}(r\geq\bm{a}^{\top}\bm{p}^{*}\mid\bm{a})\right|$
[…]
    holds for all
    $\bm{p}\in\mathcal{V}_{\bm{p}}\coloneqq\left\{\bm{p}\in\mathbb{R}^{m}:\bm{p}\geq\bm{0},\lVert\bm{p}\rVert\leq\frac{\bar{r}}{\underline{d}}\right\}$
    and $\bm{d}\in\mathcal{V}_{\bm{d}}\coloneqq[\underline{d},\bar{d}]^{m}$.
@@
    The optimal $\bm{p}^{*}$ satisfies $p_{i}^{*}=0$ if and only if $d_{i}-\mathbb{E}[a_{i}\mathbb{I}(r>\bm{a}^{\top}\bm{p}^{*})]>0$ for all $\bm{d}\in\mathcal{V}_{\bm{d}}$ and $i\in[m]$.
In this assumption, part (b) ensures that the cumulative distribution of the reward given the resource consumption request $r|\bm{a}$ is continuous and exhibits a stable growth rate. Part (c) requires strict complementarity for the optimal solutions of the stochastic program in (2), which is a non-degeneracy condition for both the primal and dual LPs.
@@
$I_{B}$ $=\{i:d_{i}-\mathbb{E}[a_{i}\mathbb{I}(r>\bm{a}^{\top}\bm{p}^{*})]=0\},$
$I_{N}$ $=\{i:d_{i}-\mathbb{E}[a_{i}\mathbb{I}(r>\bm{a}^{\top}\bm{p}^{*})]>0\}.$ (11)
By ˜2.2(c), these sets are complements, as $I_{B}\cap I_{N}=\emptyset$ and $I_{B}\cup I_{N}=\{1,...,m\}$.
3 Parallel Multi-Phase OLP Algorithm
We now present our algorithm for parallel multi-phase online learning and decision-making. Our approach is motivated by the challenges of existing methods. Specifically, the LP-based method has a strong worst-case regret bound of $\mathcal{O}(\log T)$, but its high computational costs lead to decision-making delays. Meanwhile, the first-order method updates decisions efficiently but suffers a high regret of $\mathcal{O}(\sqrt{T})$. By combining the strengths of these two methods, our new framework balances decision-making quality and computational efficiency.
@@ 3.1 Algorithm Design
Figure 1: Algorithm 1 illustration of parallel paths and the interactions between online learning and decision-making. Decisions are generated based on 1) the LP-based method (blue) with frequency $f$, 2) the first-order method (red) for the initial and final phases (with a warm start), and 3) employing the latest dual price (yellow) during intermediate phases.
@@
    We maintain two parallel paths of online learning and online decision-making. The online learning results are periodically sent to the decision-making path as a re-start point to guide subsequent updates.
@@
    In the online learning path, we employ a streamlined LP-based method, which only re-solves the updated OLP problem according to a predefined frequency, reducing the computational overhead.
@@
    In the online decision-making path, we apply the first-order method during the initial and final batches. The learning rate is optimally tuned for these two intervals.
Figure 1 illustrates our framework, which generates two parallel sequences of dual prices: $\{\bm{p}_{t}^{D}\}_{t=1}^{T}$ from the first-order method and $\{\bm{p}_{t}^{L}\}_{t=1}^{T}$ from the LP-based method. The algorithm proceeds in batches of length $f$. In the first batch, it uses the first-order method to iteratively update the dual prices, thereby guiding decision-making. At the end of this batch, it applies the LP-based method to obtain a refined dual price and passes it to the decision-making path. The LP re-solving occurs only once per batch, making the approach computationally efficient. Formally, we re-solve the OLP at every time $t$ satisfying $t\leq kf$ and $t\bmod f=0,$ where $k=\lfloor T/f\rfloor$ is the number of batches.
During intermediate batches, decisions are made based on the most recent dual prices computed from the LP at the end of the previous batch. The algorithm restarts the subgradient updates only in the final batch to guide the remaining decisions. Algorithm 1 summarizes this approach.
Algorithm 1 Parallel Multi-Phase OLP Algorithm
Input: total resource $\bm{b}$, time horizon $T$, average resource $\bm{d}=\bm{b}/T$, initial dual price $\bm{p}_{1}=\bm{0}$, re-solving frequency $f$, and number of batches $k=\lfloor T/f\rfloor$
for $t=$ 11 to $T$ do
   Observe $(r_{t},\bm{a}_{t})$ and make decision $x_{t}$ as rule (6) Update constraint for i=1,,mi=1,...,m:   remaining resource $b_{it}=b_{i,t-1}-a_{it}x_{t}$   average remaining resource $d_{it}=\frac{b_{it}}{T-t}$ if $t\leq f$ or $t\geq kf$ then
      Update learning rate $\alpha_{t}=\begin{cases}\mathcal{O}(1/f^{1/2})&t\leq f\\ \mathcal{O}(1/f^{2/3})&t\geq kf\end{cases}$ Compute subgradient and update dual price $\bm{p}_{t+1}$:
$\bm{p}_{t+1}$ $=(\bm{p}_{t}-\alpha_{t}(\bm{d}-\bm{a}_{t}x_{t}))^{+}$
@@
      Solve OLP and update dual price $\bm{p}_{t+1}$:
$\bm{p}_{t+1}=\operatorname*{argmin}_{\bm{p}\geq\bm{0}}\ \bm{d}_{t}^{\top}\bm{p}+\tfrac{1}{t}\textstyle\sum_{j=1}^{t}(r_{j}-\bm{a}_{j}^{\top}\bm{p})^{+}$
@@
Algorithm 1 balances online learning and efficient decision-making. It is responsive to dynamic environments, adapts to the latest information, and reduces computational cost by periodic re-solving. We thus establish a “wait-less” online decision-making framework where each customer’s order is processed immediately without delays from earlier requests or large-scale LP computations.
@@ 3.2 Algorithm Analysis
We decompose the performance metric of Algorithm 1 into three key components. All proofs (including essential properties of the dual price $\bm{p}_{t}$) are in Appendix B and C.
Theorem 3.1 (Performance Metric).
Under Assumptions 2.1 and 2.2, the performance $\Delta_{T}$ of Algorithm 1 satisfies
$\Delta_{T}\$$\leq\ \mu\bar{a}^{2}\textstyle\sum_{t=1}^{T}\mathbb{E}\left[\lVert\bm{p}_{t}-\bm{p}^{*}\rVert^{2}\right]+\mathbb{E}\left[\lVert(\bm{Ax}-\bm{b})^{+}\rVert\right]$
$\quad+\lVert\bm{p}^{*}\rVert\cdot\mathbb{E}\left[\lVert(\bm{b}-\bm{Ax})^{B+}\rVert\right].$ (12)
where $(\cdot)^{B+}$ indicates the projection of binding terms onto the positive orthant.
Based on the definition of $\Delta_{T}$ (10), we derive the performance metric as the sum of three key components: dual convergence, remaining resources, and binding constraint violation. This bound demonstrates that our algorithm encourages 1) smooth and balanced resource utilization, and 2) full resource consumption by the end of the time horizon. Building on Theorem 3.1, we provide a “spectrum theorem” for the algorithm’s performance.
Theorem 3.2.
Under Assumptions 2.1 and 2.2, the worst-case performance $\Delta_{T}$ of Algorithm 1 is bounded by
$\Delta_{T}\in\mathcal{O}\big(\log\left(\tfrac{T}{f}\right)+\sqrt{f}\big)$ (13)
@@
Remark 3.3 (Spectrum Theorem).
We elucidate the trade-offs in total regret induced by using LP-based methods with $\mathcal{O}(\log T)$ regret and first-order methods with $\mathcal{O}(\sqrt{T})$ regret. Our algorithm introduces a re-solving frequency $f\in[1,T]$ and achieves $\mathcal{O}(\log(T/f)+\sqrt{f})$ regret, which recovers previous results with extreme cases of $f=1$ and $f=T$. Specifically, the first-order method gives us the regret of $\sqrt{f}$ for the first batch and $f^{1/3}$ for the last batch, and the LP-based method contributes the regret of $\log(T/f)$.
Remark 3.4 (Warm Start).
We can obtain a tighter regret bound of $\mathcal{O}(\log(T/f)+f^{1/3})$ given a warm start of the initial dual price satisfying $\lVert\bm{p}_{0}-\bm{p}^{*}\rVert\leq f^{-1/3}$ or if we use the LP-based method at each time step for the first batch.
Remark 3.5 (Learning Rate Selection).
We select the best learning rate to minimize the regret upper bound consisting of (8) and (9). For the first batch, the regret grows linearly with $\alpha_{t}$ while the constraint violation is inversely proportional to $\alpha_{t}$. To achieve the tightest bound, we balance this trade-off by selecting the learning rate that minimizes the overall expression, which yields the optimal choice $\alpha_{t}=1/\sqrt{f}$ as derived in Theorem B.7. A similar analysis for the final batch leads to the choice $\alpha_{t}=1/f^{2/3}$ as shown in Theorem B.9.
@@
Theorem 3.1 decomposes the performance metrics to make the regret analysis compatible between the LP-based and first-order methods.
  • LP-based Method: As Theorem 3.1 suggests, achieving small regret requires $\bm{Ax}-\bm{b}$ to be close to zero. To enforce this, we impose a stronger condition to track the average remaining resource—if $\bm{d}_{t}$ exceeds the allowed limit, we manually set the dual prices to zero to accept all subsequent orders. This approach enables us to eliminate the “stopping time” argument and make the analysis compatible with our new framework.
  • First-order Method: The regret improvement comes from the final batch since it starts with the LP-derived learning result $\bm{p}_{kf}$, which lies within a $\mathcal{O}(1/\sqrt{kf})$-sized neighborhood around $\bm{p}^{*}$. Using the subgradient method, we bound the three components in ˜3.1 in terms of $\lVert\bm{p}_{t}-\bm{p}^{*}\rVert$ and express them in terms of $f,k,T,$ and $\alpha_{t}$. Optimizing the learning rate to $\alpha_{t}=f^{-2/3}$ yields an improved regret of $\mathcal{O}(f^{1/3})$ in the final batch.